In a two-digit number, the sum of the digits is 8. The difference between this number and the number with the digits reversed is 54. What is the number?
Answer by mananth(16946)


In a two-digit number, the sum of the digits is 8. The difference between this number and the number with the digits reversed is 54. What is the number?
Let ten's digit be x
and unit's digit be y
the value of number is 10x+y
x+y=8-----------------------(1)
10x+y-(10y+x)= 54
9x-9y =54
/9
x-y = 6-----------------(2)
add (1) and (2)
2x = 14
x=7
y=1
Number is 71